LHC 2012 - latest news

Week 17: Machine coordinators: Bernhard Holzer, Gianluigi Arduini

Plan for the week: Recover from technical stop

Plan for the restart (latest)




Link to report from morning meeting

Monday 30 April

  •  BPM checks

  • test ramp for LHCb neg. leveling plane

 8:30 meeting

Sunday 29 April

  • Start re-commissioning tests on RQF.A12

  • 14:00h Hardware checks for RQF.A12 done, restart the machine
  • 15:00h prepare injection
  • 17:30-21:30 Machine protection checks (BLMs) and tune and orbit feedback chacks
  • 21:30-03:00 Pilot ramp/squeeze: tune, chromaticity, coupling correction
  • check ADT
  • tests for the FMCM

 09:00 meeting

Saturday 28 April

  • 00:35 Trip of triplet L5 during squeeze (Power converter). Collimator functions OK during squeeze and collision process
  • 02:00 Access for RBXWTV.L2. Found cooling water valves closed.
  • 02:40 Access finished
  • Several problems (Injection handshake, functions for RCDs, BLM sanity checks sequence, SIS interlocks)
  • 06:30 staring injection and transfer line steering.
  • 09:30 Steering completed
  • 09:30-12:04 TDI checks
  • 12:04: Trip of RQF.A12. Earth fault.
  • 12:04-01:30: Investigation of earth fault on RQF.A12. Short circuit between water cooled cable and metallic box

 8:30 meeting

Friday 27 April

  • Technical stop
  • 18:30 Machine closed. End of patrols
  • Problem with the Beam Imminent Warning
  • Problem with RAMSES in the injectors and LHC requiring accesses in the injectors and LHC. ATLAS toroid trip and leak on a Xe gas line
  • 22:40 Access finished. Still problem with BIW. Start a dry ramp in the mean time to verify machine and collimator functions .
  • Temperature interlock on RBXWTV.L2

Tentative plan presented at the 08:30 meeting

Thursday 26 April

  • Technical stop


Wednesday 25 April

  • Technical stop


Tuesday 24 April

  • Technical stop


Monday 23 April

  • 01:00 Start ADT MD
  • 06:00 Start of TS

8:30 meeting

MD1 report